OpenStudy (anonymous):

What is the Least Common Denominator (LCD) of eleven-fourteenths and sixteen twenty-firsts? 7 fourteen thirty-two forty-two

OpenStudy (anonymous):

\[14=2\times 7\\ 21=3\times 7\] lcm is \(2\times 3\times 7\)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw